ali walks 120m east 150m south 40m east find the distance and displacement
1
Expert's answer
2020-01-20T05:22:53-0500

The distance


"d=d_1+d_2+d_3=120+150+40=310\:\rm m"

The displacement


"{\bf s}={\bf d}_1+{\bf d}_2+{\bf d}_3"

where


"{\bf d}_1=120{\bf i}+0{\bf j},\; {\bf d}_2=0{\bf i}-150{\bf j}, \; {\bf d}_3=40{\bf i}+0{\bf j}"

Hence


"{\bf s}=120{\bf i}+0{\bf j}+0{\bf i}-150{\bf j}+40{\bf i}+0{\bf j}=160{\bf i}-150{\bf j}"

The magnitude of a displacement


"s=\sqrt{(160)^2+(-150)^2}=219\:\rm m"